### v1

# DiMoDif: Discourse Modality-information Differentiation for Audio-visual Deepfake Detection and Localization

Abstract
Deepfake technology has rapidly advanced, posing significant threats to information integrity and societal trust. While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts but still altering the meaning, presents a more challenging detection scenario. We present a novel audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information. Our framework leverages features from deep networks that specialize in video and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery. To this end, DiMoDif employs a Transformer encoder-based architecture with a feature pyramid scheme and local attention, and optimizes the detection model through a composite loss function accounting for frame-level detections and fake intervals localization. DiMoDif outperforms the state-of-the-art on the Temporal Forgery Localization task by +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF. On the Deepfake Detection task, it outperforms the state-of-the-art by +30.5% AUC on AV-Deepfake1M, +2.8% AUC on FakeAVCeleb, and performs on-par on LAV-DF. Code available at https://github.com/mever-team/dimodif.
1 Introduction
Audio-visual deepfakes are AI-generated content involving manipulations to one or both modalities, sometimes at small time intervals, with the intention to deceive [45]. Deepfake content can widely spread online, and mislead viewers, contributing to the growing issue of information disorder. A common feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect manually. This necessitates the development of robust AI-based deepfake detection methods.
Deepfake detection is a growing area of research [54, 51, 46]. Existing approaches primarily focus on pixel-level analysis, which examines inconsistencies in gradient variations [65], color variations [30], or artifacts introduced during the generation process [11]. Also, feature-based methods analyze facial landmarks [72], temporal coherence [27], and other visual cues to identify inconsistencies in deepfake content [40]. Furthermore, there has been growing interest in detecting and localizing deepfakes in audio-visual content, with both tasks being particularly challenging due to the complex interplay between the two modalities. Key approaches include audio-visual synchronization that analyzes the consistency between visual and auditory cues namely face movements and speech [26, 9, 14], feature reconstruction learning with cross-reconstruction similarity estimation to detect anomalies prevalent in fake videos [76], and self-supervised approaches relying on the synchronization patterns learned from real videos only [28, 18].
Humans integrate multimodal sensory information, such as visual and auditory input, to extract meaningful features and perform a wide range of recognition tasks. Cognitive neuroscience research has extensively documented the interplay between video and audio [63, 64], often resulting in the alteration of perceptual content [70, 13] or even the induction of perceptual illusions [52, 60]. In the context of speech perception, the brain actively constructs content representations by combining visual, auditory, and contextual cues to predict ongoing and future utterances [12]. When visual information conflicts with auditory input, as in the McGurk effect [49] or in the case of deepfakes, increased prediction errors are observed, accompanied by significant changes in brain activity, which often manifest as higher-frequency neural oscillations and localized activation patterns, indicative of heightened cognitive effort [4].
(a) Example of speech recognition differences between lip-reading and speech-to-text models.
(b) Aggr. FakeAVCeleb [34].
(c) Aggr. LAV-DF [10].
(d) Aggr. AV-Deepfake1M [8].
Figure 1: The process of identifying audio-visual inconsistency for deepfake detection. In (a), a video’s visual and audio components are separately processed by lip-reading and speech-to-text models, then a difference score is calculated. In (b,c,d) the difference score distributions are illustrated in the form of violin plots for three different datasets.
We illustrate how state-of-the-art AI analysis tools could emulate this process in Figure 1. In Figure 1(a), an input video is first decomposed into its constituent visual and audio streams. These are then fed into state-of-the-art models for lip-reading and speech-recognition [44], and the resulting textual outputs are compared [^1] Python’s $\mathtt{difflib}$ is used (cf. https://docs.python.org/3/library/difflib.html).. Their difference is computed using the normalized Levenshtein distance, i.e. the ratio between the edit distance $\mathtt{delta}$ of the two sentences and the sum of their lengths, which results in scores between 0% (identical) and 100% (completely different). Figures 1(b), 1(c) and 1(d) illustrate aggregate results through violin plots computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. On average, lower difference is observed in real videos compared to fake ones across all datasets, indicating the suitability of this score as a measure of incongruity between audio and video. Even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which render naive thresholding on this score unsuitable for robust detection. This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, while samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
Motivated by the above, we propose Discourse Modality-information Differentiation (DiMoDif), a deep learning-based deepfake detection architecture that exploits the inter-modality differences in machine perception of speech. DiMoDif first decomposes the input video into its visual and audio streams, and extracts the respective video and audio representations based on state-of-the-art lip reading and speech-to-text pre-trained models [44]. At its core, DiMoDif employs a Transformer encoder architecture [66] with a feature pyramid scheme and local attention configuration, to spot frame-level cross-modal incongruities between the two modality-specific representations, and in that way detect and localize the deepfake forgeries. Note that low difference scores frequently occur in partly manipulated fake videos (e.g. Figure 1(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level. In addition, high difference scores occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf. Section 5.2). The architecture is optimized using a composite loss function that accounts for frame-specific detections, overlaps between predicted and ground truth fake intervals, and divergence in corresponding boundaries. We evaluate DiMoDif on Deepfake detection (DFD) and Temporal Forgery Localization (TFL) tasks using three audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. DiMoDif outperforms the state-of-the-art on FakeAVCeleb by +2.8% AUC, exhibits a significant performance increase of +30.5% AUC and +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF dataset. We also provide an analysis of DiMoDif’s generalization abilities, and an extensive ablation analysis to ground our methodological choices.
2 Related Work
2.1 Deepfake Detection Datasets
During the past years, several datasets have been proposed for video deepfake detection. Most of them focus solely on visual manipulations, thus including only the visual modality in their samples, e.g., DF-TIMIT [35], FaceForensics++ [58], DeeperForensics [32], Celeb-DF [41], WildDeepfake [81], and DF-Platter [48]. Our focus is on the detection of audio-visual deepfakes, which require content, manipulations and annotations in both modalities. Table 1 presents deepfake detection datasets that provide both visual and audio streams. DFDC [20] is a seminal large-scale benchmark in the field that contains approximately 128K samples. Although it provides audio-visual content and contains manipulations on both modalities, (i) it does not provide separate manipulation annotations per modality, (ii) fake audio parts appear in only 4% of the total number of videos, while fake visual parts in 84% [^2] Based on a later estimation: https://www.kaggle.com/datasets/basharallabadi/dfdc-video-audio-labels., and (iii) it contains videos where the depicted individual does not talk or the speaking person is not depicted and the depicted one does not talk. These features render DFDC inappropriate for audio-visual deepfake detection. KoDF [36] is another large-scale benchmark containing 237K videos of Korean subjects. Although it provides both the visual and audio components of its samples, it contains only visual manipulations, which makes it unsuitable for our problem. FakeAVCeleb [34], contains 21K manipulated and 0.5K real videos with manipulations on both modalities and corresponding annotations. Due to its size it is commonly used as an evaluation benchmark; here, we consider it for training and evaluation purposes with real video additions from the VoxCeleb2 dataset [15]. LAV-DF [10] is a recently introduced dataset that contains 136K real and fake videos with manipulations and targets on both modalities for two tasks, Deepfake Detection (DFD) and Temporal Forgery Localization (TFL). Finally, AV-Deepfake1M [8] is the largest audio-visual deepfake detection benchmark to date containing over 1.1M real and fake videos with manipulations and targets on both modalities, for both the DFD and TFL tasks. In our experimental analysis, we consider FakeAVCeleb, LAV-DF, and AV-Deepfake1M datasets, which meet all requirements of the audio-visual deepfake detection task. We also consider DFDC and KoDF for cross-dataset generalization evaluation experiments.
content manip. target
dataset size A V A V A V joint task
DFDC [20] 128,154 DFD
KoDF [36] 237,942 DFD
FakeAVCeleb [34] 21,544 DFD
LAV-DF [10] 136,304 DFD,TFL
AV-Deepfake1M [8] 1,146,760 DFD,TFL
Table 1: Deepfake detection datasets providing both video and audio. Size determines the number of samples, content specifies the provided modalities, manipulation specifies which modality has been manipulated, target specifies for which modality the dataset provides ground truth, while task is reported in the last column.
2.2 Deepfake Detection Approaches
Many deepfake detection approaches consider only video-level manipulations and processing [62, 58, 78, 38, 1, 17, 55]. Although these are only partly related to DiMoDif  we compare our method with them following previous practice. Recent works in deepfake detection have focused on leveraging both visual and audio information through learning mechanisms based on multimodal fusion, contrastive learning, and self-supervised-learning, aiming to uncover cross-modal inconsistencies in deepfake videos. Specifically, [80] demonstrated the effectiveness of an attention-based detection framework that identifies the violation of cross-modal synchronization patterns in deepfake videos, while [71] proposed a framework exploiting audio-visual inconsistency by encoding spatio-temporal information, fusing multimodal features, and then classifying based on inter- and intra-modal disharmony levels. Similarly, [56, 31, 82] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms. UMMAFormer [76] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives. The contrastive learning approach is also adopted by other works for comparing inter-modal representations within a video [14] or between different videos [77]. Other papers focus on feature similarity to identify audio-visual inconsistencies [26], and [59] compares features of the video’s lip sequence with a lip sequence generated from the audio input. Self-supervised approaches that leverage intrinsic synchronization of real videos have also been utilized [75, 25, 28, 22, 18, 53, 73, 21] to tackle the generalization challenge faced by methods trained on specific deepfake generation models. A similar motivation with ours is presented in [7, 39], which rely on a simple threshold on the divergence between visual and audio speech recognition model outputs. However, as shown in Figure 1 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M. In contrast, our approach extracts lip-reading and speech-to-text content representations that are then processed by a Transformer-based fusion mechanism that is capable of identifying discriminative speech features on top of speech divergence ones. Pre-trained lip-reading models have also been utilized as feature extractors in [29], but for visual-only deepfake detection. In addition, [57] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model. Also, specialized representation learning has been explored by [47] that considers modality emotion consistency and [69] that considers articulatory and lip movement consistency. Finally, several works [42, 50, 61, 74, 5] of the Temporal Action Localization (TAL) domain are considered as relevant by the deepfake detection literature.
3 Methodology
3.1 Problem Formulation
Consider a dataset $\mathfrak{D}=\{\mathfrak{v}_{i},y_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ where $\mathfrak{v}_{i}=\{v,a\}$ denotes a video, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^3] Utterances mainly comprise monophonic signals, besides conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples. $y_{i}\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}_{i}=\{y^{v}_{i,1},\dots,y^{v}_{i,f},y^{a}_{i,1},\dots,y^{a}_{i,f}\}$, with $y^{m}_{i,j}=(d^{m,s}_{i,j},d^{m,e}_{i,j},a^{m}_{i,j})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{i,j}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $j$, along with the corresponding distances $d^{m,s}_{i,j},d^{m,e}_{i,j}\in\mathbb{R}$ between the frame $j$ and the onset $s^{m}_{i,j}$ and offset $e^{m}_{i,j}$ of the fake interval, if the frame $j$ is fake. We train deep networks $\mathcal{F}$, the architecture of which is analysed in Section 3.2, on the DFD task $\hat{y}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ and the TFL task $\hat{\mathbf{y}}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ separately, while the predictions $\hat{\mathbf{y}}_{i}$ are decoded by:
$\hat{a}^{m}_{i,j}>0.5\text{, }\hat{s}^{m}_{i,j}=j-\hat{d}^{m,s}_{i,j}\text{, }\hat{e}^{m}_{i,j}=j-\hat{d}^{m,e}_{i,j}$ (1)
$\hat{\cdot}$ denotes prediction while its absence ground truth. Henceforth, we omit sample index $i$ to reduce notation complexity.
3.2 Architecture
Our architecture is illustrated in Figure 2. From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the lip-reading and speech-to-text models proposed in [44]. Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected. For that reason, the proposed architecture that identifies such patterns is named Discourse Modality-information Differentiation (DiMoDif).
The discourse features, $\mathbf{v}$ and $\mathbf{a}$, are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers. Then, we add trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ and employ a separation token $\mathbf{s}\in\mathbb{R}^{d}$ to encode modality information. An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{j}\in\mathbb{R}^{d}$, with $j=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [66]. $e$ depends on the task, being equal to 2 in DFD experiments and equal to 1 in TFL experiments where $\mathbf{c}$ is absent. Then, we extract $d$-dimensional feature pyramids corresponding to all $2\cdot f$ visual and audio tokens deriving from all $l$ layers of the Transformer encoder, denoted by $\mathbb{R}^{(2\cdot f)\times l\times d}\ni\mathbf{z}\subset\mathcal{T}(\mathbf{t})$. To do so, $\mathcal{T}$ processes $\mathbf{t}$ as:
$\tilde{\mathbf{z}}_{j}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{j-1}))$ (2)
$\mathbf{z}_{j}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{j}))+\tilde{\mathbf{z}}_{j}$ (3)
where $j=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [66] with $r$ number of heads, LN denotes Layer Normalization [37], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$. Then, $\mathbf{z}=\oplus\{\mathbf{z}_{j}\}_{j=1}^{l}$, omitting the tokens that correspond to $\mathbf{c}$ and $\mathbf{s}$. Given that short-term inter-modality inconsistency detection is key to our tasks, local attention is preferable in contrast to global; besides, it is computationally more efficient. Thus, we limit the Transformer’s attention to a local window $[j-q/2,j+q/2]$ of $q$ tokens around each token $j$.
Figure 2: The DiMoDif architecture.
Finally, in DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\mathbf{z}$, denoted by $\mathbf{z}^{\prime}\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
$\mathbf{z}^{\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime}\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
$\mathbf{z}^{\prime\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime\prime}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
$\hat{y}$ $=\mathbf{z}^{\prime\prime\prime}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
[…]
3.3 Objective Function
For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective. For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals. Detailed mathematical expressions explaining the computation of both loss functions can be found in the Appendix A.
4 Experimental Setup
4.1 Datasets
We consider FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8] datasets for training and evaluation. Appendix B provides details on class and split sizes. FakeAVCeleb contains 21K fake and 500 real samples, thus for balance we added 20K more from VoxCeleb2 dataset [15] including 10K of male and 10K of female speakers of several ethnicities talking in English [^4] Language identification conducted with https://github.com/speechbrain/speechbrain., reaching a total of 41,544 samples randomly split based on a 80-10-10 identity separation. LAV-DF has been released with standard splits of 78K training, 31K validation, and 26K test samples. AV-Deepfake1M has 746K training, 57K validation, and 343K test samples the metadata of which have not been released. We also evaluate DiMoDif on the test set of DFDC [20] and a randomly selected 6K-sized sample from KoDF [36] proportionally spanning all synthesis methods and real samples, to assess generalization.
4.2 Evaluation
For DFD, we use accuracy (ACC), Average Precision (AP), and area under ROC curve (AUC). For TFL, we use Average Precision at $p$ (AP@$p$), and Average Recall at $n$ (AR@$n$). We report metrics per dataset in accordance to previous works for comparability. On AV-Deepfake1M we evaluate through Codabench [^5] https://deepfakes1m.github.io/evaluation. Competing methods are audio-visual deepfake detectors, face manipulation detectors, and temporal action localization models, discussed in Section 2.
4.3 Implementation Details
We train DiMoDif for 100 epochs with early stopping patience 10 and checkpoint based on validation metrics, AUC for DFD and the sum of AP@$p$ and AR@$n$ for TFL. Batch size is set to 64, initial learning rate to 0.001 with ReduceLROnPlateau scheduler [^6] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html and Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$. Also, we set a maximum sequence length $f$ to 600 and zero-pad the smaller. We consider a grid for $\mathcal{T}$’s size, with $(d,r,u)\in\{(32,2,1),(64,2,1),(64,4,1),(128,4,2),(256,8,2)\}$ and number of layers $l\in\{1,3,5\}$. Grid search for AV-Deepfake1M is performed with 200K training samples to reduce resource requirements, but retrain the best configuration on the whole set. We spent $\sim$3900 GPU hours for training on NVIDIA GeForce RTX 3090 Ti GPUs.
5 Results
5.1 Temporal Forgery Localization
Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
MDS [14] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
AGT [50] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
BMN [42] $\mathcal{V}$ 24.0 7.6 0.1 53.3 41.2 31.6 26.9
BMN (I3D) [42] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
ActionFormer [74] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
UMMAFormer [76] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
Table 2: Temporal forgery localization results on LAV-DF [10]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Method Modality AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@50 AR@30 AR@20 AR@10 AR@5
Meso4 [1] $\mathcal{V}$ 09.86 06.05 02.22 00.59 38.92 38.91 38.81 36.47 26.91
MesoInception4 [1] $\mathcal{V}$ 08.50 05.16 01.89 00.50 39.27 39.22 39.00 35.78 24.59
EfficientViT [17] $\mathcal{V}$ 14.71 02.42 00.13 00.01 27.04 26.99 26.43 23.90 20.31
TriDet+VideoMAEv2 [61, 67] $\mathcal{V}$ 21.67 05.83 00.54 00.06 20.27 20.23 20.12 19.50 18.18
TriDet+InternVideo [61, 68] $\mathcal{V}$ 29.66 09.02 00.79 00.09 24.08 24.06 23.96 23.50 22.55
ActionFormer+VideoMAEv2 [74, 67] $\mathcal{V}$ 20.24 05.73 00.57 00.07 19.97 19.93 19.81 19.11 17.80
ActionFormer+InternVideo [74, 68] $\mathcal{V}$ 36.08 12.01 01.23 00.16 27.11 27.08 27.00 26.60 25.80
BA-TFD [10] $\mathcal{AV}$ 37.37 6.34 0.19 0.02 45.55 40.37 35.95 30.66 26.82
BA-TFD+ [9] $\mathcal{AV}$ 44.42 13.64 0.48 0.03 48.86 44.51 40.37 34.67 29.88
UMMAFormer [76] $\mathcal{AV}$ 51.64 28.07 7.65 1.58 44.07 43.93 43.45 42.09 40.27
DiMoDif (ours) $\mathcal{AV}$ 86.93 75.95 28.72 5.43 81.57 80.85 80.25 78.84 76.64
Table 3: Temporal forgery localization results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Tables 2 and 3 present the performance of our method in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8] datasets, respectively. On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50,20}, is second to best wrt AR@10, and third wrt AP@{0.5,0.75,0.95}. UMMAFormer [76], BA-TFD+ [9] and ActionFormer [74] exhibit similar performance, while the rest of the models have lower localization abilities. On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase. The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
Method Modality AUC
F3-Net [55] $\mathcal{V}$ 52.0
MDS [14] $\mathcal{AV}$ 82.8
EfficientViT [17] $\mathcal{V}$ 96.5
BA-TFD [10] $\mathcal{AV}$ 99.0
UMMAFormer [76] $\mathcal{AV}$ 99.8
DiMoDif (ours) $\mathcal{AV}$ 99.6
Table 4: Deepfake detection results on LAV-DF [10]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Method Modality AUC ACC
Video-LLaMA (13B) E5 [75] $\mathcal{AV}$ 50.7 25.1
LipForensics [29] $\mathcal{V}$ 51.6 68.8
Face X-Ray [38] $\mathcal{V}$ 61.5 73.8
Meso4 [1] $\mathcal{V}$ 50.2 75.0
MesoInception4 [1] $\mathcal{V}$ 50.1 75.0
SBI [62] $\mathcal{V}$ 65.8 69.0
MDS [14] $\mathcal{AV}$ 56.6 59.4
DiMoDif (ours) $\mathcal{AV}$ 96.3 83.3*
Table 5: Deepfake detection results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. *Validation set accuracy is reported as Codabench only provides AUC score.
Category
Method Modality Pre-training RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
Unsupervised AVBYOL [25] $\mathcal{AV}$ LRW [16] 50.0 50.0 73.4 61.3 88.7 80.8 60.2 33.8 73.2 61.0 73.9 59.2
VQ-GAN [21] $\mathcal{V}$ LRS2 [2] - - 50.3 49.3 57.5 53.0 49.6 48.0 62.4 56.9 55.0 51.8
AVAD [22] $\mathcal{AV}$ LRS2 [2] 62.4 71.6 93.6 93.7 95.3 95.8 94.1 94.3 93.8 94.1 94.2 94.5
AVAD [22] $\mathcal{AV}$ LRS3 [3] 70.7 80.5 91.1 93.0 91.0 92.3 91.6 92.7 91.4 93.1 91.3 92.8
Supervised Xception [58] $\mathcal{V}$ ImageNet [19] - - 88.2 88.3 92.3 93.5 67.6 68.5 91.0 91.0 84.8 85.3
LipForensics [29] $\mathcal{V}$ LRW [16] - - 97.8 97.7 99.9 99.9 61.5 68.1 98.6 98.7 89.4 91.1
AD DFD [80] $\mathcal{AV}$ Kinetics [33] 74.9 73.3 97.0 97.4 99.6 99.7 58.4 55.4 100. 100. 88.8 88.1
FTCN [78] $\mathcal{V}$ 96.2 97.4 100. 100. 77.4 78.3 95.6 96.5 92.3 93.1
RealForensics [28] $\mathcal{V}$ LRW [16] - - 88.8 93.0 99.3 99.1 99.8 99.8 93.4 96.7 95.3 97.1
DiMoDif (ours) $\mathcal{AV}$ 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
Table 6: Deepfake detection results on FakeAVCeleb [34]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. Supervised methods that use pre-training are fine-tuned on FakeAVCeleb, while unsupervised methods are not trained with labels and fake examples.
Test dataset
FakeAVCeleb [34] LAV-DF [10] AV-Deepfake1M [8] DFDC [20] KoDF [36]
ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC
Training dataset FakeAVCeleb [34] 98.1 99.9 99.9 39.3 59.5 61.9 25.3 55.1 55.1 51.4 62.7 60.8 65.5 77.4 32.7
LAV-DF [10] 43.6 70.7 71.3 95.6 99.6 99.6 32.1 69.7 66.2 51.8 52.9 54.5 43.8 88.8 66.5
AV-Deepfake1M [8] 56.3 82.0 85.4 59.2 83.2 81.5 87.7 97.3 97.3 51.3 50.5 49.7 77.0 77.7 33.4
Table 7: Generalization performance on Deepfake detection task.
AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@100 AR@50 AR@30 AR@20 AR@10 AR@5
Test dataset: LAV-DF
Training dataset LAV-DF [10] 95.5 87.9 54.6 20.6 94.2 93.7 93.2 92.7 91.4 89.2
AV-Deepfake1M [8] 59.8 26.2 2.3 0.2 75.3 72.3 69.7 67.3 62.5 56.2
Test dataset: AV-Deepfake1M
LAV-DF [10] 16.8 4.6 0.3 0.0 31.9 30.1 28.6 27.0 23.6 19.9
AV-Deepfake1M [8] 96.0 86.2 34.5 6.9 89.2 88.7 88.2 87.7 86.4 84.3
Table 8: Generalization performance on Temporal Forgery Localization task.
5.2 Deepfake Detection
Tables 5, 5 and 6, present comparative results in terms of video-level Deepfake Detection (DFD) on LAV-DF [10], AV-Deepfake1M [8], and FakeAVCeleb [34], respectively. On LAV-DF, DiMoDif performs on-par with the state-of-the-art method UMMAFormer [76], achieving 99.6 AUC, better than the remaining competitive methods. On AV-Deepfake1M dataset, DiMoDif outperforms all state-of-the-art competitive methods by a significant absolute performance increase of +30.5% in terms of AUC [^7] Accuracy (ACC) is also reported by other papers; however, Codabench does not provide it for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase.. On FakeAVCeleb, DiMoDif outperforms all competitive methods on average achieving +4.1% absolute AP increase, and +2.8% absolute AUC increase, as well as in most forgery types. Finally, challenging real samples with normalized Levenshtein distance over 0.8 (cf. Section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt a naive distance-thresholding approach.
5.3 Generalization
This paper is the first to conduct generalization experiments on the used datasets. Table 7 presents DiMoDif’s cross-dataset performance on DFD. Training on FakeAVCeleb [34] leads to poor performance on LAV-DF, AV-Deepfake1M, and DFDC but to moderate performance on KoDF with 77.4 AP. Training on LAV-DF [10] leads to moderate performance on FakeAVCeleb and AV-Deepfake1M, low performance on DFDC but good performance on KoDF with 88.8 AP. Training on AV-Deepfake1M [8] dataset leads to high scores on FakeAVCeleb and LAV-DF with 85.4 AUC and 81.5 AUC respectively, moderate performance on KoDF with 77.7 AP, but low performance on DFDC. In general, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.1. Our models focus on cross-modal information differences, while many DFDC videos do not contain talking or the talking person is different from the depicted one; thus, a visual-only face manipulation detector would likely lead to better results. Appendix C provides detailed generalization evaluation on FakeAVCeleb with scores per forgery type. Table 8 presents cross-dataset performance on TFL. Training on AV-Deepfake1M leads to acceptable performance on LAV-DF, in contrast to training on LAV-DF and evaluating on AV-Deepfake1M.
5.4 Ablation Analysis
(a) Window size $q$ (LAV-DF)
(b) Window size $q$ (AV-Deepfake1M)
(c) Feature pyramid use (LAV-DF)
(d) Feature pyramid use (AV-Deepfake1M)
(e) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
(f) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
Figure 3: Ablation and hyperparameter tuning analysis.
Figure 3 presents an ablation analysis that supports our methodological choices. For AV-Deepfake1M we report validation performance as conducting the ablations through Codabench would be impractical. These results pertain to TFL, while further ablations as well as the same analysis wrt DFD can be found in Appendix D. Figures 3(a) and 3(b) indicate that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$. Also, the feature pyramid scheme provides a small performance increase as indicated by Figures 3(c) and 3(d). The size of $\mathcal{T}$ plays crucial role in achieving maximum performance as shown in Figures 3(e) and 3(f), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$. Specifically, the larger the model the higher its performance.
6 Conclusions
In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech. It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities. It considers a video and audio speech recognition feature extraction stage, and a Transformer encoder-based learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries. An extensive evaluation study indicates the effectiveness of our framework reaching state-of-the-art performance and in several cases surpassing it. Our model is limited by the languages recognized by the used speech-recognition backbones, although it exhibits good generalization to Korean. Also, it assumes one continually visible speaking person, at least moderate video quality and face size to detect face landmarks, and requires both modalities for manipulation detection. Future work includes joint training on Deepfake Detection and Temporal Forgery Localization tasks to assess potential benefits and extension to multiple speakers.
7 Acknowledgments
This work has been partially funded by the Horizon Europe projects AI4Trust (GA No. 101070190) and vera.ai (GA No. 101070093).
Appendix A Objective function details
For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective:
$\mathfrak{L}_{DFD}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{j=1}^{l}y[m]\text{log}\hat{y}[j,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[j,m])$ (7)
where $j$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
For the TFL task, we consider a combination of three loss functions. For clarity we omit to also include the symbols for the averaging across $l$ layers, which however takes place as well. Specifically, we use:
  1. 1.
    The focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent among each video’s frames:
    $\mathfrak{L}_{focal}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}-\alpha_{t}^{m,j}(1-p_{t}^{m,j})^{\gamma}\text{log}(p_{t}^{m,j})$ (8)
    in which:
    $\alpha_{t}^{m,j}=\begin{cases}\alpha&\text{if }y^{m}_{j}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
    $p_{t}^{m,j}=\begin{cases}\hat{y}^{m}_{j}&\text{if }y^{m}_{j}=1\\ 1-\hat{y}^{m}_{j}&\text{otherwise}\end{cases}$ (10)
    while $\alpha$ and $\gamma$ are hyperparameters.
  2. 2.
    The DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals:
    $\mathfrak{L}_{DIoU}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{j},\mathbf{b}^{m}_{j})+\\ \frac{\rho^{2}(\hat{b}^{m}_{j},b^{m}_{j})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,j}$ (11)
    where $j$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{j}=[s^{m}_{j},e^{m}_{j}]$ denotes a time interval, $b^{m}_{j}=0.5\cdot(e^{m}_{j}+s^{m}_{j})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake.
  3. 3.
    The smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals:
    $\mathfrak{L}_{1}^{smooth}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{j}-\hat{s}^{m}_{j})\\ +h(e^{m}_{j}-\hat{e}^{m}_{j})\bigg{)}\cdot\mathds{1}_{m,j}$ (12)
    in which $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake, and $h$ is defined as:
    $h(x)=\begin{cases}0.5x^{2}&\text{if }\lvert x\rvert<1\\ \lvert x\rvert-0.5&\text{otherwise}\end{cases}$ (13)
We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\mathds{1}_{m,j}$:
$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{focal}+\mathfrak{L}_{DIoU}+\mathfrak{L}_{1}^{smooth}\big{)}/p$ (14)
We finally average both $\mathfrak{L}_{DFD}$ and $\mathfrak{L}_{TFL}$ across batch samples during training.
Appendix B Dataset class and split sizes
Table 9 provides details on dataset class and split sizes. Although, FakeAVCeleb paper [34] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples instead of the commonly reported 20,000 size for this dataset. Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”. Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
FakeAVCeleb LAV-DF AV-Deepfake1M
train val test total train val test total train val test total
FVFA 8,678 1,052 1,105 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
FVRA 7,760 975 974 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
RVFA 399 49 52 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
RVRA 17,766 1,535 1,199 20,500 21,254 8,271 6,906 36,431 186,666 14,235 - -
total 34,603 3,611 3,330 41,544 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
Table 9: Dataset class and split sizes. F: fake, R: real, V: video, A: audio.
Appendix C Generalization on FakeAVCeleb
Table 10 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M especially in some cases such as FVFA-WL and FVFA-FS reaching 95.6 AUC and 94.5 AUC, respectively. Training on LAV-DF also results in good generalization especially in FVFA-FS case with 90.0 AUC.
Category
Training dataset RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
FakeAVCeleb [34] 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
LAV-DF [10] 40.2 74.3 77.1 87.2 78.2 87.8 72.1 90.0 70.0 86.9 74.4 88.0
AV-Deepfake1M [8] 18.8 80.8 63.5 80.3 85.6 95.6 74.2 94.5 75.9 93.8 74.8 91.1
Table 10: Detailed generalization performance for each forgery type of the FakeAVCeleb dataset.
Appendix D Ablation Analysis (further experiments)
Figures 4(a) and 4(b) indicate little to no sensitivity to the focal loss’ hyperparameter $\alpha$. Figures 4(c) and 4(d) indicate that the use of a learning rate scheduler is beneficial however the type of scheduler is not significantly affecting the results.
(a) Focal loss hyperparmeter $\alpha$ (LAV-DF)
(b) Focal loss hyperparmeter $\alpha$ (AV-Deepfake1M)
(c) Learning rate scheduler (LAV-DF)
(d) Learning rate scheduler (AV-Deepfake1M)
Figure 4: Ablation analysis on Temporal Forgery Localization task.
Figure 5 indicates that on DFD window size $q=15$ is again optimal, but performance is not increased with feature pyramids nor with different learning rate schedules. Also, medium-sized models achieve maximum performance.
(a) Window size $q$ (FakeAVCeleb)
(b) Window size $q$ (LAV-DF)
(c) Window size $q$ (AV-Deepfake1M)
(d) Feature pyramid use (FakeAVCeleb)
(e) Feature pyramid use (LAV-DF)
(f) Feature pyramid use (AV-Deepfake1M)
(g) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (FakeAVCeleb)
(h) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
(i) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
(j) Learning rate scheduler (FakeAVCeleb)
(k) Learning rate scheduler (LAV-DF)
(l) Learning rate scheduler (AV-Deepfake1M)
Figure 5: Ablation and hyperparameter tuning analysis on Deepfake Detection task.